# Resilience Against Soft Faults through Adaptivity in Spectral Deferred Correction

### v1

@@ 1 Introduction
Resilience against faults can be achieved via hardware and software. Dedicated software resilience strategies often come with significant overhead, which exacerbates the computational efficiency that can be gained by adaptive methods in comparison. The potential for adaptive time stepping schemes to improve resilience as well as computational efficiency has already been explored for embedded Runge-Kutta methods (RKMs) [9] as well as spectral deferred correction (SDC) with adaptive iteration number [10].
SDC is a time stepping scheme that iteratively solves fully implicit RKMs. It can be advantageous over the more common diagonally implicit RKMs for stiff or complex problems [11, 12], when employing complex splitting techniques [13, 14, 15] or as a basis for parallel-in-time (PinT) methods [16, 17, 18, 19]. PinT [20] codes especially require attention to resilience because they are designed to run on large portions of modern HPC systems. After adaptive step-size selection for SDC has been shown to boost computational efficiency [21], we now demonstrate its capacity to provide resilience against soft faults on par with a dedicated resilience strategy. We do this by performing numerical experiments where we manually insert faults and compute recovery rates for different resilience strategies.
@@ 2.1 Spectral Deferred Correction
where $u(t)$ is the solution at time $t$, the subscript $t$ marks a derivative with respect to time and $f(u)$ describes the temporal evolution of $u$. Note that we restrict this discussion to autonomous scalar problems for the sake of simpler notation, but the method can be applied to partial differential equations (PDEs) and non-autonomous problems as well.
@@
where $\tilde{q}_{mj}$ is the preconditioner. The matrix containing the $\tilde{q}_{mj}$ is called $Q_{\Delta}$ in the literature and is usually chosen to be lower triangular so that it can be inverted by forward substitution.
@@ 2.2 Adaptive step size selection in SDC
The collocation problem corresponds to the equation for the stages of a fully implicit RKM. SDC with fixed number of iterations can also be interpreted as an RKM itself. This suggests that it should be possible to adopt established methods from general RKM to the special case of SDC. It has been shown that efficient step size adaptive algorithms can be constructed for SDC along the lines of embedded RKM [21].
The two key ingredients are an estimate of the local error for the current step size, and an estimate of the optimal step size that would have resulted in the target local error. In embedded RKM [26], the error is estimated by computing two solutions of orders $p$ and $q>p$ and subtracting them to get an estimate of the lower order method
@@
In SDC, an error estimate for fixed number of iterations can be constructed from the increment, as the order increases by one with each iteration. The resulting scheme is called $\Delta t$-adaptivity [21]. It is also possible to choose both step size and iteration number adaptively in a scheme called $\Delta t$-$k$-adaptivity [21]. Iteration number is chosen by declaring convergence of the collocation problem based on the residual or the increment. Then, the error for step size adaptivity is estimated by comparing the order $M$ solution at the second to last collocation node $u_{M-1}$ to an order $M-1$ interpolation to $t=\tau_{M-1}$ of the polynomial defined by $\{u_{i}:i=0,1,...,M-2,M\}$.
@@ 2.3 Resilience
The potential of adaptive SDC to improve resilience depends on its ability to detect and recover from faults. The SDC residual should indicate soft faults that alter the solution. The range of faults that SDC can recover from, after detecting them by increased residual, will depend on the preconditioner and the problem. The Newton solver in a non-linear problem, for instance, may not converge from a faulty initial guess. Similarly, convergence is not always guaranteed in SDC. Still, for a wide range of faults, $k$-adaptive SDC has been shown to provide an adequate solution in the presence of faults [10]. We will call SDC with fixed step size and adaptive iteration number “$k$-adaptivity” in the remainder of this article.
When choosing the step size adaptively ($\Delta t$-adaptivity), we expect to detect the fault in the estimate of the local error. Recovery happens by rejection of steps with local error exceeding the target. This is already an effective resilience strategy for embedded RKM [9]. $\Delta t$-$k$-adaptivity, in turn, can detect faults in the residual as well as in the local error estimate and can recover by means of restart if SDC does not converge after continued iteration. Note that the tolerance determining the accuracy of adaptive SDC is a natural tolerance for fault detection. If the fault is insignificant, it is not necessary to expend resources on recovery. Adaptive SDC will only cause additional iterations or restarts if the fault perturbs the solution above the general accuracy requirement.
@@
To set a baseline, we compare against the dedicated resilience strategy Hot Rod [29], which was developed for explicit RKM, but can be adapted to SDC. It is a detector for soft faults that can be tuned to high accuracy and faults can be corrected by restarting. Detection works by computing two estimates of the local error and taking the difference to get an even more accurate estimate that is very sensitive to faults. For the first estimate, we use the increment in SDC. The second estimate computes a secondary solution at the current time by extrapolating from previous times using a Taylor method and comparing to the solution obtained by SDC. The original publication includes a Taylor method with sufficient order for Cash-Karp’s method. We compute Taylor methods at runtime in order to facilitate the easy tuning of order that is one of the advantages of SDC.
Note that both error estimates need to compute the local error of the lower order method. This means we have to advance in time also with the lower order method and use the final iteration only for computing the error. This is in contrast to step step size adaptive schemes, which compute the local error of the lower order method, but advance with the most accurate solution available. For good performance it is necessary to select a detection threshold such that significant faults will be corrected, but insignificant faults or false positives do not cause too much overhead. In the original paper, thresholds for different levels of resilience requirements are obtained by means of machine learning. In our experiments, we simply select a problem specific tolerance by (human) trial and error.
@@
A fault can be inserted in a variety of ways. We can insert in any iteration, at any collocation node, in any space position, bit, and at any time. We pick a single time for fault insertion up front in order to limit the number of options. We then insert faults at random in up to 4000 experiments for the nonlinear Schrödinger, Allen-Cahn and Rayleigh-Benard example For the Lorenz attractor problem, there are only 3840 combinations for inserting faults at a fixed time because there are 5 options for iteration number, 4 choices for collocation node (including initial conditions), 3 solution components, each with 64 bits.
Since we want the same faults for all strategies, we restrict the random insertion to faults that would target the fixed strategy in the same way. That is, when we perform a fixed number of $k_{\mathrm{max}}$ iterations in the fixed and $\Delta t$-adaptivity strategies, we insert faults only in the first $k_{\mathrm{max}}$ iterations, even if the $k$- and $\Delta t$-$k$-adaptivity strategies may perform more iterations in this step.
@@
To decide if a fault is recovered, we compute the global error and accept it as recovered if it is no greater than 10 % larger compared to the global error in a fault-free run. Since the solution is not arbitrarily accurate in any case, slightly larger error is usually acceptable. The recovery rate is then the ratio of experiments where the fault was recovered to the total number of experiments.
3 Benchmark Problems
We now introduce the four test problems that we use in numerical experiments. All of them are non-linear. One is an ordinary differential equation and three are partial differential equations. We denote the simulation time at which we insert faults as $t_{\mathrm{fault}}$. All implementations are publicly available on GitHub [^2] See https://github.com/Parallel-in-Time/pySDC for the repository as part of the pySDC library [31] and continuously tested. Numerical experiments are run on the JUSUF supercomputer at Forschungszentrum Jülich [32].
3.1 Lorenz attractor
@@
$x_{t}$ $=\sigma\left(y-x\right),$ (6)
$y_{t}$ $=\rho x-y-xz,$ (7)
$z_{t}$ $=xy-\beta z,$ (8)
@@
$\sigma$ $=10,~{}\rho=28,~{}\beta=8/3,$
$x$ $(t=0)=y(t=0)=z(t=0)=1,$
$t$ $\in[0,20],$
$t_{\mathrm{fault}}$ $=10,$
@@
We solve the problem implicitly with a self-constructed Newton solver and use the SciPy [35] method solve_ivp from the integrate package with an explicit embedded RKM of order 5(4) [36] with tolerances close to machine precision to obtain reference solutions.
3.2 Non-linear Schrödinger
@@
with periodic boundary conditions. The global error is computed with respect to the analytic solution [38, Equation (39)]. The solution is shown in Figure 2.
Figure 2: Initial condition for the Schrödinger example (left) and the solution at the end of the simulation at $t=1$ (right). The initial condition is purely real, but the solution shifts to the complex domain as the simulation progresses.
3.3 Allen-Cahn
@@
Figure 3: Left: Initial condition consisting of a circle of high phase embedded in the low phase. Right: Evolution of the radius of the circle over time computed from the analytical solution. The simulation runs until $t=0.025$.
@@
3.4 Rayleigh-Benard Convection
@@
with $u$ denoting the velocity in $x$-direction, $v$ the velocity in $z$-direction, $T$ the temperature, $p$ the pressure, and $\kappa=\sqrt{Ra\cdot Pr}$ and $\nu=\sqrt{Ra/Pr}$, where $Ra=2\times 10^{4}$ and $Pr=1$ are the Rayleigh and Prandtl numbers. We choose a spatial domain of size $\Omega=[0,8)\times(-1,+1)$, with $256\times 128$ degrees of freedom, periodic boundary conditions in $x$-direction, Dirichlet boundary conditions
@@
in $z$-direction and $\int_{\Omega}p=0$ for the pressure gauge. We discretize the problem with a pseudo-spectral method using a Fourier base horizontally and an ultraspherical base [39] vertically. We treat the linear parts, which are written on the left hand side of Equation 23, implicitly in spectral space and the non-linear convective terms on the right hand side explicitly in physical space.
@@
4 Numerical Results
We test the resilience strategies discussed in Section 2.3 for the four benchmarks described above. We compare to SDC with fixed iteration number and step size as a reference with no special resilience capabilities, which we call the “fixed” scheme. Before presenting how the strategies perform for all problems from Section 3, we investigate them in detail with the Lorenz attractor problem. Figure 5 and Figure 6 illustrate for two specific experiments how susceptible this problem is to small perturbations.
Figure 7: Recovery rate for the Lorenz attractor problem depending on the acceptance threshold. The threshold is shown as a ratio to the fault-free error, meaning a value of one requires fault correction to deliver exactly the same error as a simulation without faults, A value of two would count twice the error as without faults as recovered. Any threshold larger than one leads to high recovery rates for all recovery strategies. We choose a threshold of 1.1 for computing recovery rates.
In practice, values between one and two are probably reasonable. We find that in the case of the chaotic Lorenz attractor the exact value within this range does not matter much.
Figure 8 shows recovery rates dependent on the collocation node that is targeted by the fault
Note that in the pySDC implementation, collocation node 0 stores the initial conditions. This means faults that hit in collocation node 0 cannot be recovered by restarting, since the initial conditions will continue to include the fault. This causes significantly lower recovery rate for faults in node 0 than in other nodes. We also find that the fixed strategy has slightly lower recovery rate in node 3 than in nodes 1 and 2. After SDC has converged, the solution to the step is computed as a weighted sum of the $u_{m}$ during the so called collocation update. With the Gauß-Radau quadrature rule with three nodes that we use here, the last node is the end point of the interval and the collocation update amounts to simply copying $u_{3}$. Hence, when a fault targets any but the last node in the last iteration, it has no impact on the solution to the step.
Since faults in the last iteration and any but the last node do not affect the collocation update, the fixed and $\Delta t$-adaptivity strategies have the highest recovery rate in the last iteration. Because the fault nevertheless impacts the residual, strategies with adaptive $k$ may continue to iterate and propagate the fault into the collocation update. Keep in mind that Hot Rod performs one extra iteration, but advances in time with the solution of the fifth iteration. If a fault occurs in the initial conditions in the fifth iteration, Hot Rod will trigger a restart and the fault ends up affecting the solution. Otherwise, the recovery rate of the fixed strategy decreases with iteration number. The earlier the fault occurs, the more iterations are available to smoothe out the perturbation.
The resilience properties we identified for the Lorenz problem carry over to the PDE examples from Section 3 as well. We show the recovery rate for recoverable faults only in Figure 12.
Figure 12: Comparison of recovery rate for theoretically recoverable faults for all problems from Section 3. The Schrödinger and Rayleigh-Benard problems use complex128 numbers instead of the usual float64, which we show as two float64 numbers back to back. The impact of faults appears to be the same per bit regardless of whether the complex or the real part is targeted.
5 Discussion
Acknowledgements
We thankfully acknowledge funding from the European High-Performance Computing Joint Undertaking (JU) under grant agreement No 955701 and 101118139. The JU receives support from the European Union’s Horizon 2020 research and innovation programme and Belgium, France, Germany, and Switzerland. We also thankfully acknowledge funding from the German Federal Ministry of Education and Research (BMBF) grant 16HPC048.
The authors gratefully acknowledge computing time granted on JUSUF through the CSTMA project at Jülich Supercomputing Centre.
Availability of supporting data
For instructions on how to reproduce the results shown in this paper, please consult the project section in the documentation of the pySDC code at https://parallel-in-time.org/pySDC/projects/Resilience.html.

### v2

@@ 1 Introduction
Resilience against faults can be achieved via hardware and software. Dedicated software resilience strategies often come with significant overhead, which exacerbates the computational efficiency that can be gained by adaptive methods in comparison. The potential for adaptive time stepping schemes to improve resilience as well as computational efficiency has already been explored for embedded Runge-Kutta methods (RKMs) [5] as well as spectral deferred correction (SDC) with adaptive iteration number [19]. Adaptive methods generally measure accuracy and trigger additional work if the accuracy was found to be insufficient. While these mechanisms are primarily developed to react to the behaviour of the problem, some transient soft faults can be corrected in adaptive methods after they have been detected in the accuracy measurement.
SDC is a time stepping scheme that iteratively solves the equations for the stages of fully implicit RKMs with preconditioned Picard iterations. It can be advantageous in terms of computational efficiency over the more common diagonally implicit RKMs for stiff or complex problems [22, 15], when employing complex splitting techniques [34, 28, 31] or as a basis for parallel-in-time (PinT) methods [14, 36, 8, 16]. PinT [17] codes especially require attention to resilience because they are designed to run on large portions of modern HPC systems. After adaptive step-size selection for SDC has been shown to boost computational efficiency [4], we now demonstrate its capacity to provide resilience against soft faults on par with a dedicated resilience strategy. We do this by performing numerical experiments where we manually insert faults and compute recovery rates for different resilience strategies.
@@ 2.1 Spectral Deferred Correction
where $u(t)$ is the solution at time $t$, the subscript $(\cdot)_{t}$ marks a derivative with respect to time and $f(u)$ describes the temporal evolution of $u$. Note that we restrict this discussion to autonomous scalar problems for the sake of simpler notation, but the method can be applied to partial differential equations (PDEs) and non-autonomous problems as well.
@@
where $\tilde{q}_{mj}$ is the preconditioner. The matrix containing the $\tilde{q}_{mj}$ is called $Q_{\Delta}$ in the literature and is lower triangular, allowing inversion by forward substitution.
@@ 2.2 Adaptive step size selection in SDC
The collocation problem that is solved in SDC can be viewed as a fully implicit RKM. SDC with fixed number of iterations can also be interpreted as an RKM itself. This suggests that it should be possible to adopt established methods from general RKM to the special case of SDC. It has been shown that efficient step size adaptive algorithms can be constructed for SDC along the lines of embedded RKM [4].
The two key ingredients are an estimate of the local error for the current step size, and an estimate of the optimal step size that would have resulted in the target local error. In embedded RKM [24], the error is estimated by computing two solutions of orders $p$ and $q>p$ and subtracting them to get an estimate of the error of the lower order method
@@
In SDC, an error estimate for fixed number of iterations can be constructed from the increment, as the order increases by one with each iteration. The resulting scheme is called $\Delta t$-adaptivity [4]. It is also possible to choose both step size and iteration number adaptively in a scheme called $\Delta t$-$k$-adaptivity [4]. The iteration number is chosen by stopping when the residual of the collocation problem or the increment reach sufficiently small values. Then, the error for step size adaptivity is estimated by comparing the order $M$ solution at the second to last collocation node $u_{M-1}$ to an order $M-1$ interpolation to $t=\tau_{M-1}$ of the polynomial defined by $\{u_{i}:i=0,1,...,M-2,M\}$.
@@ 2.3 Resilience
The potential of adaptive SDC to improve resilience depends on its ability to detect and recover from faults. The residual of the collocation problem should be increased by soft faults that alter the solution. The range of faults that SDC can recover from, after detecting them by increased residual, will depend on the preconditioner and the problem. A plain Newton solver in a non-linear problem, for instance, may not converge from a faulty initial guess. Similarly, convergence is not always guaranteed in SDC. Still, for a wide range of faults, $k$-adaptive SDC has been shown to provide an adequate solution in the presence of faults by performing additional SDC iterations [19]. We will call SDC with fixed step size and adaptive iteration number “$k$-adaptivity” in the remainder of this article.
When choosing the step size adaptively ($\Delta t$-adaptivity), we expect to detect the fault in the estimate of the local error. Recovery happens by rejection of steps with local error exceeding the target. This is already an effective resilience strategy for embedded RKM [5]. $\Delta t$-$k$-adaptivity, in turn, can detect faults using the residual as well as the local error estimate and can recover by means of restart if SDC does not converge after continued iteration. Note that the tolerance determining the accuracy of adaptive SDC is a natural tolerance for fault detection. If the fault is insignificant, it is not necessary to expend resources on recovery. Adaptive SDC will only cause additional iterations or restarts if the fault perturbs the solution above the general accuracy requirement.
@@
To set a baseline, we compare against the dedicated resilience strategy Hot Rod [21], which was developed for explicit RKM, but can be adapted to SDC. It is a detector for soft faults that can be tuned to high accuracy and faults can be corrected by restarting. Detection works by computing two estimates $\epsilon_{1}$ and $\epsilon_{2}$ of the local error and taking the difference to get a quantity $\Delta$ that is very sensitive to faults
$\Delta=\left|\epsilon_{1}-\epsilon_{2}\right|_{\infty}=\left|\delta^{(p)}+\mathcal{O}(\Delta t^{q})-\delta^{(p)}+\mathcal{O}(\Delta t^{r})\right|_{\infty}=\mathcal{O}(\Delta t^{p+2}),$ (6)
with $q,r>p+1$. For $\epsilon_{1}$, we use the increment in SDC. For $\epsilon_{2}$, we compute a secondary solution at the current time via a linear multistep method (LMM) [24, Chapter III.1], which uses the solution values from previous timesteps, and compare to the solution obtained by SDC, taking the accumulation of local errors within the LMM into account. The original publication of Hot Rod [21] includes the error estimate [7, Equation 1.3] with sufficient order for the RKM Cash-Karp’s method [10]:
$\begin{split}\epsilon_{2}^{CK}&=\frac{1}{30}\left(u_{n-3}+18u_{n-2}-9u_{n-1}-10u_{n}\right)+\frac{\Delta t}{10}\left(3f_{n-2}+6f_{n-1}+f_{n}\right)\\ &=\delta^{(4)}+\mathcal{O}(\Delta t^{6}),\end{split}$ (7)
where $u_{n}$ is the solution to the current timestep, $u_{n-i}$ is the solution to the timestep $i$ steps prior, and $f_{n-i}=f(u_{n-i})$. We compute the coefficients of error estimates for $\epsilon_{2}$ as in Equation 7 at runtime in order to facilitate the easy tuning of order that is one of the advantages of SDC.
Since the increment is an estimate of the error of the second to last iterate in SDC, that is $\epsilon_{1}=\delta^{(k_{\mathrm{max}}-1)}$ with $k_{\mathrm{max}}$ SDC iterations, $\epsilon_{2}$ has to also estimate the error of the second to last iterate in order to satisfy Equation 6. This means we have to advance in time also with the lower order method and use the final iteration only for computing $\epsilon_{1}$. This is in contrast to step step size adaptive schemes, which compute the local error of the lower order method, but advance with the most accurate solution available. For good performance it is necessary to select a detection threshold such that significant faults will be corrected, but insignificant faults or false positives do not cause too much overhead. In the original Hot Rod paper [21], thresholds for different levels of resilience requirements are obtained by means of machine learning. In our experiments, we simply select a problem specific tolerance by (human) trial and error.
@@
A fault can be inserted in a variety of ways. We can insert in any iteration, at any collocation node, in any space position, bit, and at any time. We pick a single time for fault insertion up front in order to limit the number of options. All strategies that we test are fifth order accurate, meaning $k_{\mathrm{max}}=5$ for strategies with fixed iteration number, except $k_{\mathrm{max}}=6$ for Hot Rod. Since we want the same faults for all strategies, we restrict the random insertion to the first five iterations. We then insert faults at random in up to 4000 experiments for PDE examples. We also test one ordinary differential equation example, for which there are only 3840 combinations for inserting faults in the solution variables at a fixed time because there are 5 options for iteration number, 4 choices for collocation node (including initial conditions), 3 solution components, each with 64 bits.
@@
To decide if the simulation has recovered from the fault, we compute the global error at the end of the simulation and accept successful recovery if it is not increased beyond some relative threshold compared to a fault-free run. Since the solution is not arbitrarily accurate in any case, slightly larger error is usually acceptable. The recovery rate is then the ratio of experiments where the fault was recovered to the total number of experiments.
2.4 Benchmark Problems
We employ four non-linear test problems in numerical experiments, which we describe in detail in Appendix A. We use the Lorenz attractor (Appendix A.1) as a toy problem that is very sensitive to faults and allows to study the resilience properties in detail. Then, we test the resilience capabilities for challenging PDEs, discretized in space with spectral discretizations. The PDEs are the non-linear Schrödinger equation (Appendix A.2), the Allen-Cahn reaction-diffusion equation (Appendix A.3), and the incompressible fluid dynamics benchmark Rayleigh-Benard convection (Appendix A.4). All implementations are publicly available on GitHub [^2] See https://github.com/Parallel-in-Time/pySDC for the repository as part of the pySDC library [37] and continuously tested. Numerical experiments are run on the JUSUF supercomputer at Forschungszentrum Jülich [40].
3 Numerical Results
We test the resilience strategies discussed in Section 2.3 for the four benchmarks described in Appendix A. We compare to SDC with fixed iteration number and step size as a reference with no special resilience capabilities, which we call the “fixed” scheme. Before presenting how the strategies perform for the PDE benchmarks, we investigate them in detail with the Lorenz attractor problem. Figure 1 and Figure 2 illustrate for two specific experiments how susceptible this problem is to small perturbations.
Figure 3: Recovery rate for the Lorenz attractor problem depending on the acceptance threshold. The threshold is shown as the ratio of the global error to the global error of a fault-free run, meaning a value of one requires fault correction to deliver exactly the same error as a simulation without faults. A value of two would count twice the error as without faults as recovered. Any threshold larger than one leads to high recovery rates for all recovery strategies. We choose a threshold of 1.1 for computing recovery rates.
We find that in the case of the chaotic Lorenz attractor the exact value within this range does not matter much. Note that the acceptable size of perturbations due to faults is dependent on the application and the value of 10 % may not be suitable for all users.
Figure 4 shows recovery rates dependent on the collocation node that is targeted by the fault.
Note that in the pySDC implementation, collocation node 0 stores the initial conditions. This means faults that occur in collocation node 0 cannot be recovered by restarting, since the initial conditions will continue to include the fault. This causes significantly lower recovery rate for faults in node 0 than in other nodes. We also find that the fixed strategy has slightly lower recovery rate in node 3 than in nodes 1 and 2. After SDC has converged, the solution to the step is computed as a weighted sum of the $u_{m}$ during the so called collocation update. With the Gauß-Radau quadrature rule with three nodes that we use here, the last node is the end point of the interval and the collocation update amounts to simply copying $u_{3}$. Hence, when a fault targets any but the last node in the last iteration, it has no impact on the solution to the step.
Since faults in the last iteration and any but the last node do not affect the collocation update, the fixed and $\Delta t$-adaptivity strategies have the highest recovery rate in the last iteration. Because the fault nevertheless impacts the residual, strategies with adaptive $k$ may continue to iterate and propagate the fault into the collocation update. Keep in mind that Hot Rod performs one extra iteration, but advances in time with the solution of the fifth iteration. If a fault occurs in the initial conditions in the fifth iteration, Hot Rod will trigger a restart and the fault ends up affecting the solution. Otherwise, the recovery rate of the fixed strategy decreases with iteration number. The earlier the fault occurs, the more iterations are available to smooth out the perturbation.
The resilience properties we identified for the Lorenz problem carry over to the PDE examples from Appendix A as well. We show the recovery rate for recoverable faults only in Figure 8.
Figure 8: Comparison of recovery rate for theoretically recoverable faults for all problems from Section 2.4. The Schrödinger and Rayleigh-Benard problems use complex128 numbers instead of the usual float64, which we show as two float64 numbers back to back. The impact of faults appears to be the same per bit regardless of whether the complex or the real part is targeted.
4 Discussion
Acknowledgements
We thankfully acknowledge funding from the European High-Performance Computing Joint Undertaking (JU) under grant agreement No 955701 and 101118139. The JU receives support from the European Union’s Horizon 2020 research and innovation programme and Belgium, France, Germany, and Switzerland. We also thankfully acknowledge funding from the German Federal Ministry of Education and Research (BMBF) grant 16HPC048. The authors gratefully acknowledge computing time granted on JUSUF through the CSTMA project at Jülich Supercomputing Centre. The authors gratefully acknowledge the helpful comments from the reviewers.
Availability of supporting data
For instructions on how to reproduce the results shown in this paper, please refer to the instructions in pySDC/projects/resilience/README.rst in version 5.6 of pySDC [38].
Appendix A Benchmark Problems
Here, we describe the benchmark problems used in numerical experiments in detail. We denote the simulation time at which we insert faults as $t_{\mathrm{fault}}$.
A.1 Lorenz attractor
@@
$\begin{split}x_{t}&=\sigma\left(y-x\right),\\ y_{t}&=\rho x-y-xz,\\ z_{t}&=xy-\beta z,\end{split}$ (8)
@@
$\begin{split}\sigma&=10,~{}\rho=28,~{}\beta=8/3,\\ x&(t=0)=y(t=0)=z(t=0)=1,\\ t&\in[0,20],\\ t_{\mathrm{fault}}&=10,\end{split}$ (9)
@@
We integrate Equation 8 implicitly in the SDC iterations using a self-constructed Newton solver and use the SciPy [39] method solve_ivp from the integrate package with an explicit embedded RKM of order 5(4) [11] with tolerances close to machine precision to obtain reference solutions.
A.2 Non-linear Schrödinger
@@
with periodic boundary conditions. The global error is computed with respect to the analytic solution [3, Equation (39)]. The solution is shown in Figure 10.
Figure 10: Initial condition for the Schrödinger example (left) and the solution at the end of the simulation at $t=1$ (right). The initial condition is purely real, but the solution shifts to the complex domain as the simulation progresses. Reproduced from [4] under CC BY 4.0 licence.
A.3 Allen-Cahn
@@
Figure 11: Left: Initial condition consisting of a circle of high phase embedded in the low phase. Right: Evolution of the radius of the circle over time computed from the analytical solution. The simulation runs until $t=0.025$. Reproduced from [4] under CC BY 4.0 licence.
@@
A.4 Rayleigh-Benard Convection
@@
with $u$ denoting the velocity in $x$-direction, $v$ the velocity in $z$-direction, $T$ the temperature, $p$ the pressure, $\kappa$ the thermal diffusivity, and $\nu$ the kinematic viscosity. We use a Rayleigh number of
$Ra=\frac{\Delta TL_{z}^{3}}{\kappa\nu}=3.2\times 10^{5},$ (25)
with $\Delta T=2$ the temperature difference between the top and bottom plates and $L_{z}=2$ the separation between the top and bottom plates, and a Prandtl number of
$Pr=\frac{\nu}{\kappa}=1.$ (26)
We choose a spatial domain of size $\Omega=[0,8)\times(-1,+1)$, with $256\times 128$ degrees of freedom, periodic boundary conditions in $x$-direction, Dirichlet boundary conditions
@@
in $z$-direction and $\int_{\Omega}p=0$ for the pressure gauge. We discretize the problem with a pseudo-spectral method using a Fourier base in $x$-direction and an ultraspherical base [32] in $z$-direction. We treat the linear parts, which are written on the left hand side of Equation 24, implicitly in spectral space and the non-linear convective terms on the right hand side explicitly in physical space.